To throw the discus, the thrower holds it with a fully outstretched arm. Starting from rest, he begins to turn with a constant angular acceleration, releasing the discus after making one complete revolution. The diameter of the circle in which the discus moves is about 1.7 m . If the thrower takes 1.1 s to complete one revolution, starting from rest, what will be the speed of the discus at release?
Answer #1

radius is r = d/2 = 1.7/2 = 0.85 m

time period is T = 1.1sec
initial velocity is vo = 0 m/s


speed of the discus is v = (2*pi*R)/T = (2*3.142*0.85)/1.1 = 4.86 m/sec